OpenStudy (anonymous):

Which is not a factor of 56? A. 4 B. 8 C. 14 D. 26

OpenStudy (solomonzelman):

\(\Large\color{black}{ \bf 56=2 \times 2 \times 2 \times 7 }\) \(\Large\color{black}{ \bf 26=\color{red} { 13 } \times 2 }\) \(\Large\color{black}{ \bf 14=2 \times 7 }\) \(\Large\color{black}{ \bf 8=2 \times 2 \times 2 }\) \(\Large\color{black}{ \bf 4=2 \times 2 }\)

OpenStudy (anonymous):

The factors for 56 are all numbers between −56 and 56, which divides 56 evenly. Check numbers between −56 and 56 Since 56 divided by −56 is a whole number, −56 is a factor for 56. −56 Since 56 divided by −28 is a whole number, −28 is a factor for 56. −28 Since 56 divided by −14 is a whole number, −14 is a factor for 56. −14 Since 56 divided by −8 is a whole number, −8 is a factor for 56. −8 Since 56 divided by −7 is a whole number, −7 is a factor for 56. −7 Since 56 divided by −4 is a whole number, −4 is a factor for 56. −4 Since 56 divided by −2 is a whole number, −2 is a factor for 56. −2 Since 56 divided by −1 is a whole number, −1 is a factor for 56. −1 Since 56 divided by 1 is a whole number, 1 is a factor for 56. 1 Since 56 divided by 2 is a whole number, 2 is a factor for 56. 2 Since 56 divided by 4 is a whole number, 4 is a factor for 56. 4 Since 56 divided by 7 is a whole number, 7 is a factor for 56. 7 Since 56 divided by 8 is a whole number, 8 is a factor for 56. 8 Since 56 divided by 14 is a whole number, 14 is a factor for 56. 14 Since 56 divided by 28 is a whole number, 28 is a factor for 56. 28 Since 56 divided by 56 is a whole number, 56 is a factor for 56. 56 The factors for 56 are −56,−28,−14,−8,−7,−4,−2,−1,1,2,4,7,8,14,28,56. −56,−28,−14,−8,−7,−4,−2,−1,1,2,4,7,8,14,28,56

OpenStudy (nikosis01):

d